### v1

# Asymptotics for partitions over the Fibonacci numbers and related sequences

Abstract.
In this paper, harkening back to ideas of Hardy and Ramanujan, Mahler and de Bruijn, with the addition of more recent results on the Fibonacci Dirichlet series, we determine the asymptotic number of ways $p_{F}(n)$ to write an integer as the sum of non-distinct Fibonacci numbers, as well as an asymptotic for the summatory function of the numbers $q_{F}(n)$—the number of ways to write $n$ as the sum of distinct Fibonacci numbers. This appears to be the first such asymptotic result concerning non-distinct partitions over Fibonacci numbers. As well, under weak conditions, we prove analogous results for a general linear recurrences.
Key words and phrases:
Partitions, recurrence sequences, Dirichlet series
2010 Mathematics Subject Classification:
Primary 11P82; Secondary 11M41
1. Introduction
We consider two functions each of which describes different types of partitions over Fibonacci numbers, $F_{k}$. In particular, for a positive integer $n$, write
(1) $n=a_{2}F_{2}+a_{3}F_{3}+\cdots+a_{k}F_{k}+\cdots.$
We let $p_{F}(n)$ denote the number of non-distinct partitions of $n$ over the Fibonacci sequence, that is, solutions to (1) in nonnegative integers $a_{k}$, and let $q_{F}(n)$ denote the number of distinct partitions of $n$ over the Fibonacci sequence, that is, solutions to (1) with $a_{k}\in\{0,1\}$. In recent work, Chow and Slattery [2] gave results on $q_{F}(n)$ (denoted $R(n)$ in their paper), and noted that their work shows that neither $q_{F}(n)$, nor its partial sums, have a ‘nice’ asymptotic formula. In particular, they showed that there is some oscillation in the partial sums of $q_{F}(n)$, and gave bounds on these oscillations. In this paper, we determine asymptotic formulas for both $p_{F}(n)$ and the partial sums of $q_{F}(n)$, and describe these oscillations.
The asymptotic theory of partitions goes back to the celebrated result Hardy and Ramanujan who, in 1918, showed that the number of ways $p(n)$ to write a positive integer as the sum of positive integers satisfies $p(n)\sim(4n\sqrt{3})^{-1}e^{\pi\sqrt{2n/3}},$ as $n\to\infty$. Here, $p(n)$ has an asymptotic with a non-oscillating main term. Mahler [6] and de Bruijn [1] encountered a partition asymptotic with an oscillatory main term—they considered the number of ways $p_{r}(n)$ of writing $n$ as the sum of non-distinct $r$th powers, for a positive integer $r\geqslant 2$. Here, we contribute the following result.
Theorem 1.
Let $p_{F}(n)$ and $q_{F}(n)$ be the number of partitions of $n$ over non-distinct and distinct Fibonacci numbers, respectively. Then, as $n\to\infty$,
$p_{F}(n)=\psi_{p_{F}}\left(\frac{\log n}{\log\varphi}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\varphi}+\frac{\log\log\varphi}{\log\varphi}+\frac{\log 5-\log\varphi+4\gamma}{2\log\varphi}-2\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
and,
$\sum_{m\leqslant n}q_{F}(m)=\psi_{q_{F}}\left(\frac{\log n}{\log\varphi}\right)n^{\frac{\log 2}{\log\varphi}}\left(1+O\left(\frac{1}{n}\right)\right),$
where $\psi_{p_{F}}(x)$ and $\psi_{q_{F}}(x)$ are explicit positive $1$-periodic functions, $\varphi$ is the golden mean and $\gamma$ is the Euler–Mascheroni constant.
Theorem 1 seems to be the first asymptotic result chracterising non-distinct partitions over Fibonacci numbers. Distinct partitions over Fibonacci numbers have received considerable attention; see [2] and the references therein for details. In very recent work, addressing a question of Chow and Slattery [2], Kempton [5] has shown that $n^{-\frac{\log 2}{\log\varphi}}\sum_{m\leqslant n}q_{F}(m)$ is $\log$-periodic. Our proof of Theorem 1 gives an explicit description of this $\log$-periodic function.
This paper is organised as follows. In Sections 2 and 3, we focus on the case of $p_{F}(n)$ of Theorem 1. In particular, inspired by ideas of de Bruijn [1], we prove an exact formula for the generating function of $p_{F}(n)$ in Section 2. We then use a saddle-point method in Section 3 to give the asymptotic result for $p_{F}(n)$. In Section 4, we prove our main result for $q_{F}(n)$; this is done a bit more quickly and with less detail than the previous two sections since the methods and results are similar. Finally, in Section 5, we give the complete extension of Theorem 1 to the case of a linear recurrence with dominant root. In particular, suppose that $P_{k}$ is a strictly increasing linearly recurrent sequence of positive integers of degree at least $2$ with $P_{1}=1$, such that the characteristic polynomial $\chi_{P}(x)$ of $P_{k}$ has a single dominant root $\beta>0$ and $P_{k}=\lambda\beta^{k}+\lambda_{2}\beta_{2}^{k}+\cdots+\lambda_{r}\beta_{r}^{k},$ where $\lambda,\lambda_{2},\ldots,\lambda_{r}$ are constants and $\beta_{2},\ldots,\beta_{r}$ are the algebraic conjugates of $\beta$. In Section 5, we prove
Theorem 2.
Let $p_{P}(n)$ and $q_{P}(n)$ be the number of partitions of $n$ over non-distinct and distinct members of the sequence $P_{k}$, respectively. Then, as $n\to\infty$,
$p_{P}(n)=\psi_{p_{P}}\left(\frac{\log n}{\log\beta}\right)(\log n)^{-\frac{1}{2}}n^{\left(\frac{\log n}{2\log\beta}+\frac{\log\log\beta}{\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{3}{2}\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
and for any fixed $\varepsilon\in(0,1)$,
$\sum_{m\leqslant n}q_{P}(m)=\psi_{q_{P}}\left(\frac{\log n}{\log\beta}\right)n^{\frac{\log 2}{\log\beta}}\left(1+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)\right),$
where $\psi_{p_{P}}(x)$ and $\psi_{q_{P}}(x)$ are explicitly determinable positive $1$-periodic functions.
Remark 1.
Note that if $\beta>2$, then Theorem 2 implies the well-known result that the number of integers that have only ones and zeros in their base-$\beta$ expansion has density $\log 2/\log\beta<1$. This number is also the Hausdorff dimension of the subset of real numbers in the interval $[0,1]$ that have only ones and zeros in their base-$\beta$ expansions. $\Diamond$
2. An exact formula for for the generating series of $p_{F}(n)$
To prove Theorem 1, we study the asymptotics of the generating series
$F_{2}(x):=\sum_{n\geqslant 0}p_{F}(n)x^{n}=\prod_{k\geqslant 2}\left(1-x^{F_{k}}\right)^{-1}$
as $x\rightarrow 1^{-}$. Note that we are starting with $F_{2}=1$, and not with $F_{1}=1$, since we do not want two representations of $1$. We will necessarily need to consider a Fibonacci Dirichlet series. Navas [8] determined most of the properties we need by considering the analytic continuation of the series $\zeta_{F}(z)=\sum_{k\geqslant 1}F_{k}^{-z}$, but here with $1$ doubly represented. It turns out this is not much of a problem. To deal with this, we merely consider the product
(2) $F(x)=(1-x)^{-1}F_{2}(x)=\prod_{k\geqslant 1}\left(1-x^{F_{k}}\right)^{-1},$
then translate the results back to $F_{2}(x)$. It is convenient to do a change of variables, setting $x=e^{-s}$, so that we are considering the function $F(e^{-s})$, with particular interest in the asymptotics as $s\rightarrow 0^{+}$.
Taking the logarithm and using the Taylor series of the logarithm near $1$,
(3) $\log F(e^{-s})=-\sum_{k=1}^{\infty}\log\left(1-e^{-sF_{k}}\right)=\sum_{k=0}^{\infty}\sum_{m=1}^{\infty}\frac{1}{m}e^{-sF_{k}m}.$
Mellin’s formula for the exponential function states that for $a>0$ and $w>0$,
[…]
Inserting this into (3) and interchanging integration and summation, as we may by absolute convergence, and finally rearranging the sum,
[…]
where $a>0$, $\zeta(z+1)$ denotes the Riemann $\zeta$-function at $z+1$, and, as above $\zeta_{F}(z):=\sum_{k\geqslant 1}F_{k}^{-z}$. Note that $\zeta_{F}(z)$ is absolutely convergent for $\Re(z)>0$ and continuable to a meromorphic function on all of $\mathbb{C}$; this is discussed more below—see Navas [8].
We would like to estimate the final integral in (4) using Cauchy’s formula; that is, we would like to move the vertical contour towards $\infty$- ∞. Applying the functional equation of the Riemann $\zeta$-function, we find that
[…]
where $n$ runs over the positive integers, and, as we shall see in what follows, the vertical line avoids poles of the integrand. Also, contributions from horizontal paths of integration do not contribute in the limit, as they are moved up or down respectively. Thus, the integral in (4) is nothing but the sum over the residues at the poles of the integrand. In what follows, we prove
Theorem 3.
The function $F_{2}(z)$ defined above satisfies, as $s\to 0^{+}$,
$\log F_{2}(e^{-s})=\frac{(\log s)^{2}}{2\log\varphi}-(\log s)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(s)+O(s).$
where $f(s)=f(\varphi s)+O(s^{2})$, $c_{3}:=\frac{1}{2}\log 5-\frac{1}{2}\log\varphi+2\gamma,$ and
$c_{2}:=\frac{3\gamma^{2}}{2\log\varphi}-\frac{\gamma_{1}}{\log\varphi}+\frac{\pi^{2}}{12\log\varphi}+2\gamma\left(\frac{\log 5-\log\varphi}{2\log\varphi}\right)+\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}.$
The proof of Theorem 3 will come as a direct application of five propositions, each having to do with contributions coming from certain singularities of $s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)$, and one lemma. To this end, note that for $s>0$, the function $z\mapsto s^{-z}$ can be expanded in a Taylor series as
(5) $s^{-z}=1-z\log s+\tfrac{1}{2}(\log s)^{2}z^{2}+O(z^{3}),$
which converges for all $z\in\mathbb{C}$. The $\Gamma$-function has simple poles at non-positive integers, where the residue at $-n$ is
$\underset{{{s}={-n}}}{\textnormal{Res}}\left\{\Gamma(s)\right\}=\frac{(-1)^{n}}{n!}.$
The function $\zeta(z+1)$ has a simple pole with residue $1$ at $z=0$, and the trivial zeros of $\zeta(z+1)$ at all points in $-2\mathbb{N}-1$ will cancel the corresponding poles of $\Gamma$.
The function $\zeta_{F}(z)$ is the most mysterious of the functions we will consider here, but much is known due to work of Navas [8]. We will use several of these properties, which we have gathered into the following proposition.
Proposition 1 (Navas, 2001).
The Dirichlet series $\zeta_{F}(z)$ can be continued analytically to a meromorphic function on all of $\mathbb{C}$, still called $\zeta_{F}(z)$, whose singularities are simple poles at $s=s(n,k)=-2k+\tfrac{\pi i(2n+k)}{\log\varphi},$ for $n,k\in\mathbb{Z}$ and $k\geqslant 0$, with
$\underset{{{z}={s}}}{\textnormal{Res}}\left\{\zeta_{F}(z)\right\}=\frac{(-1)^{k}5^{s/2}{-s\choose k}}{\log\varphi}.$
Moreover, we have that $\zeta_{F}(-(4m+2))=0$ for each $m\in\mathbb{N}_{0}$, and $\zeta_{F}(-1)=-1$.
The poles of $\zeta_{F}(z)$ collude with the poles of $s^{-z}\Gamma(z)\zeta(1+z)$ to become poles of higher order. Combining our knowledge of the functions $s^{-z},$ $\Gamma(z)$ and $\zeta(1+z)$ with Proposition 1, the integrand $s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)$ has
  • a triple pole at $z=0$,
  • a simple pole coming from $\Gamma(z)$ at $z=-1$,
  • double poles at $z\in-4\mathbb{N}$, and
  • simple poles off the real line at $z=s(n,k)$, for $k\geqslant 0$ and $k\neq-2n$.
Note that the poles of $\Gamma(z)$ at $z=-(4m+2)$ for $m\in\mathbb{N}_{0}$ are cancelled by the corresponding zero of $\zeta_{F}(z)$. The higher order poles require more terms in the expansions of $\Gamma(z)$, $\zeta(1+z)$ and $\zeta_{F}(z)$. We will consider these in the order above.
To calculate the contribution of the triple pole at $z=0$, we need the first three terms of the Laurent expansions of the contributing functions, i.e., of $\Gamma(z)$, $\zeta(z+1)$ and $\zeta_{F}(z)$. The former two are well known,
(6) $\Gamma(z)=\frac{1}{z}+\gamma+\tfrac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)z+O(z^{2}),$
where $\gamma$ is the Euler–Mascheroni constant, and
(7) $\zeta(1+z)=\frac{1}{z}+\gamma-\gamma_{1}z+O(z^{2}),$
where $\gamma_{1}$ is the first Stieltjes constant,
$\gamma_{1}=\lim_{m\rightarrow\infty}\left(\sum_{k=1}^{m}\frac{\log k}{k}-\tfrac{1}{2}(\log m)^{2}\right).$
For $\zeta_{F}(z)$, results beyond the residue are not available in literature, so we present them here.
Lemma 1.
Near $z=0$, we have
$\zeta_{F}(z)=\frac{1}{\log\varphi}\cdot\frac{1}{z}+\frac{\log 5-\log\varphi}{2\log\varphi}+\left(\frac{\log\varphi-3\log 5}{12}+\frac{(\log 5)^{2}}{8\log\varphi}+c_{1}\right)z+O(z^{2}),$
where $c_{1}:=\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}\approx-0.20436188.$
Proof.
The first two terms were found in [8]. For clarity, we repeat the same process here. As a first step, noting $\bar{\varphi}=-1/\varphi$ and $F_{n}=(\varphi^{n}-\bar{\varphi}^{n})/\sqrt{5}$, we have
(8) $\zeta_{F}(z)=\sum_{n\geqslant 1}\frac{1}{F_{n}^{z}}=5^{z/2}\sum_{n\geqslant 1}\frac{1}{(\varphi^{n}-\bar{\varphi}^{n})^{z}}.$
We recover the Taylor expansion of $5^{z/2}$ from that of the exponential function,
(9) ${5}^{z/2}=e^{\log(5)z/2}=\sum_{k\geqslant 0}\frac{(\log 5)^{k}}{2^{k}k!}z^{k}=1+\frac{\log 5}{2}z+\frac{(\log 5)^{2}}{8}z^{2}+O(z^{3}).$
For the series, we start with an application of the binomial theorem to give
$\sum_{n\geqslant 1}\frac{1}{(\varphi^{n}-\bar{\varphi}^{n})^{z}}=\sum_{n\geqslant 1}\frac{1}{\varphi^{nz}}\left(1+(-1)^{n+1}\varphi^{-2n}\right)^{-z}=\sum_{n\geqslant 1}\frac{1}{\varphi^{nz}}\sum_{k\geqslant 0}\binom{-z}{k}(-1)^{k(n+1)}\varphi^{-2nk}.$
Since the double sum is absolutely convergent, as argued in [8], we swap the order of summation and recognise a geometric series to give
$\sum_{n\geqslant 1}\frac{1}{(\varphi^{n}-\bar{\varphi}^{n})^{z}}$ $=\sum_{k\geqslant 0}\binom{-z}{k}(-1)^{k}\sum_{n\geqslant 1}\left((-1)^{k}\varphi^{-(z+2k)}\right)^{n}$
$=\sum_{k\geqslant 0}\binom{z+k-1}{k}\frac{(-1)^{k}\varphi^{-(z+2k)}}{1-(-1)^{k}\varphi^{-(z+2k)}}$
$=\sum_{k\geqslant 0}\frac{\Gamma(z+k)}{\Gamma(z)\Gamma(k+1)}\cdot\frac{(-1)^{k}}{\varphi^{z+2k}+(-1)^{k+1}}$
$=\frac{1}{\varphi^{z}-1}+\sum_{k\geqslant 1}\frac{\Gamma(z+k)}{\Gamma(z)k!}\cdot\frac{(-1)^{k}}{\varphi^{z+2k}+(-1)^{k+1}}.$
The first term is particularly cumbersome, but a few applications of L’Hôpital’s rule gives
$\frac{1}{\varphi^{z}-1}=\frac{1}{\log\varphi}\cdot\frac{1}{z}-\frac{1}{2}+\frac{\log\varphi}{12}z+O(z^{2}).$
For the remaining summands, by $\tfrac{1}{\Gamma(z)}=z+O(z^{2})$, $\Gamma(z+k)=\Gamma(k)+O(z)=(k-1)!+O(z)$, and
$\frac{(-1)^{k}}{\varphi^{z+2k}+(-1)^{k+1}}=\frac{(-1)^{k}}{\varphi^{2k}+(-1)^{k+1}}+O(z),$
we have that
$\sum_{n\geqslant 1}\frac{1}{(\varphi^{n}-\bar{\varphi}^{n})^{z}}=\frac{1}{\log\varphi}\cdot\frac{1}{z}-\frac{1}{2}+\left(\frac{\log\varphi}{12}+c_{1}\right)z+O(z^{2}),$
where $c_{1}:=\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}\approx-0.20436188.$ Thus, using (9),
$\zeta_{F}(z)$ $=5^{z/2}\left(\frac{1}{\log\varphi}\cdot\frac{1}{z}-\frac{1}{2}+\left(\frac{\log\varphi}{12}+c_{1}\right)z+O(z^{2})\right)$
$=\frac{1}{\log\varphi}\cdot\frac{1}{z}+\left(-\frac{1}{2}+\frac{\log 5}{2\log\varphi}\right)+\left(\frac{\log\varphi}{12}+c_{1}-\frac{\log 5}{4}+\frac{(\log 5)^{2}}{8\log\varphi}\right)z+O(z^{2})$
$=\frac{1}{\log\varphi}\cdot\frac{1}{z}+\frac{\log 5-\log\varphi}{2\log\varphi}+\left(\frac{\log\varphi-3\log 5}{12}+\frac{(\log 5)^{2}}{8\log\varphi}+c_{1}\right)z+O(z^{2}),$
which is the desired result. ∎
Proposition 2.
We have
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=\frac{(\log s)^{2}}{2\log\varphi}-\left(\frac{1}{2}\log 5-\frac{1}{2}\log\varphi+2\gamma\right)\frac{\log s}{\log\varphi}+c_{2},$
where
$c_{2}:=\frac{3\gamma^{2}}{2\log\varphi}-\frac{\gamma_{1}}{\log\varphi}+\frac{\pi^{2}}{12\log\varphi}+2\gamma\left(\frac{\log 5-\log\varphi}{2\log\varphi}\right)+\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}.$
Proof.
Around $z=0$, using (5), (6) and (7), we have
$s^{-z}\Gamma(z)\zeta(1+z)=\frac{1}{z^{2}}+\frac{2\gamma-\log s}{z}+\left(\frac{(\gamma-\log s)(3\gamma-\log s)}{2}-\gamma_{1}+\frac{\pi^{2}}{12}\right)+O(z).$
Thus, by Lemma 1, we have
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=\frac{1}{\log\varphi}\left(\frac{(\gamma-\log s)(3\gamma-\log s)}{2}-\gamma_{1}+\frac{\pi^{2}}{12}\right)\\ +\frac{\log 5-\log\varphi}{2\log\varphi}(2\gamma-\log s)+\left(\frac{\log\varphi-3\log 5}{12}+\frac{(\log 5)^{2}}{8\log\varphi}+c_{1}\right)$
where $c_{1}:=\sum_{k\geqslant 1}\frac{(-1)^{k}}{k(\varphi^{2k}+(-1)^{k+1})}\approx-0.20436188.$ Gathering powers of $\log s$ finishes the proof. ∎
The simple pole at $z=-1$ is a straightforward calculation, using known values.
Proposition 3.
We have
$\underset{{{z}={-1}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=-\frac{s}{2}.$
Proof.
We calculate, using that $\underset{{{z}={-1}}}{\textnormal{Res}}\left\{\Gamma(z)\right\}=-1$, $\zeta(0)=-1/2,$ and $\zeta_{F}(-1)=-1$ to give
$\underset{{{z}={-1}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=s\,\zeta(0)\,\zeta_{F}(-1)\cdot\underset{{{z}={-1}}}{\textnormal{Res}}\left\{\Gamma(z)\right\}=s\left(\frac{-1}{2}\right)(-1)(-1)=-\frac{s}{2}.$
For the calculation of $\zeta_{F}(-1)$ see Navas [8, Eq. 9]
For the double poles at $z\in-4\mathbb{N}$, we require the first two terms of the Laurent expansions of $\Gamma(z)$ and $\zeta_{F}(z)$ respectively around these points. For $\Gamma(z)$, the first term of the expansion is well known to be $\frac{1}{(4n)!}\cdot\frac{1}{z+4n}$. The constant term is surprisingly difficult to find in literature, but it can be easily calculated as the derivative of $(z+4n)\Gamma(z)$, evaluated at $z=-4n$. In order to accomplish this, we repeatedly apply the functional equation $z\Gamma(z)=\Gamma(z+1)$ to find that
$\Gamma(z)=\frac{\Gamma(z+1)}{z}=\dots=\frac{\Gamma(z+4n+1)}{z(z+1)\cdots(z+4n)},$
so that we only need to evaluate the derivative
$\frac{d}{dz}\left\{\frac{\Gamma(z+4n+1)}{z(z+1)\cdots(z+4n-1)}\right\}$
at $z=-4n$. Recalling that $\Gamma(1)=1$ and $\Gamma^{\prime}(1)=-\gamma$, we find that the constant term of the Laurent series of $\Gamma(z)$ around $z=-4n$ is
$\frac{d}{dz}\left\{\frac{\Gamma(z+4n+1)}{z(z+1)\cdots(z+4n-1)}\right\}\Big{|}_{z=-4n}=\frac{1}{(4n)!}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right),$
so that near $z=-4n$,
(10) $\Gamma(z)=\frac{1}{z+4n}\cdot\frac{1}{(4n)!}+\frac{1}{(4n)!}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+O(z+4n).$
We also require the constant term of the Laurent series of $\zeta_{F}(z)$ around $z=-4n$. Following Navas [8],
$\zeta_{F}(z)$ $=5^{z/2}\sum_{k=0}^{\infty}\binom{-z}{k}\frac{(-1)^{k}}{\phi^{z+2k}+(-1)^{k+1}}=5^{z/2}\sum_{k=0}^{\infty}\frac{\Gamma(1-z)}{\Gamma(1-z-k)k!}\cdot\frac{(-1)^{k}}{\phi^{z+2k}+(-1)^{k+1}}$
$=5^{z/2}\frac{\Gamma(1-z)}{\Gamma(1-z-2n)(2n)!}\cdot\frac{1}{\phi^{z+4n}-1}+5^{z/2}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{\infty}\frac{\Gamma(1-z)}{\Gamma(1-z-k)k!}\cdot\frac{(-1)^{k}}{\phi^{z+2k}+(-1)^{k+1}}.$
The second term is holomorphic at $z=-4n$ and contributes to the constant term with its value, which we denote by $c_{-4n}$. For the first term, we note that
$\lim_{z\rightarrow-4n}5^{z/2}=5^{-2n},\qquad\mbox{and}\qquad\lim_{z\rightarrow-4n}\frac{\Gamma(1-z)}{\Gamma(1-z-2n)(2n)!}=\frac{(4n)!}{((2n)!)^{2}}.$
Thus, it remains to note that
$\lim_{z\rightarrow-4n}\frac{d}{dz}\left\{\frac{z+4n}{\phi^{z+4n}-1}\right\}=-\frac{1}{2},$
which is easily shown by differentiating and applying L’Hôpital’s rule twice. Proposition 1 gives that the residue of $\zeta_{F}(z)$ at $z=-4n$ is
$b_{-4n}:=\underset{{{z}={-4n}}}{\textnormal{Res}}\left\{\zeta_{F}(z)\right\}=\frac{5^{-2n}}{\log\phi}\binom{4n}{2n}=\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}},$
so that near $z=-4n$, we have
(11) $\zeta_{F}(z)=\frac{b_{-4n}}{z+4n}+\left(c_{-4n}-\frac{5^{-2n}}{2}\cdot\frac{(4n)!}{((2n)!)^{2}}\right)+O(z+4n).$
Proposition 4.
We have
$g(s):=\sum_{n=1}^{\infty}\underset{{{z}={-4n}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}=\sum_{n=1}^{\infty}\alpha_{n}\,s^{4n}-\log s\sum_{n=1}^{\infty}\beta_{n}s^{4n},$
where $\beta_{n}:={b_{-4n}\zeta(1-4n)}/{(4n)!}$,
$\alpha_{n}:=\frac{B_{2n}}{4n\cdot(4n)!}\left(b_{-4n}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+c_{-4n}-\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+\frac{b_{-4n}}{(4n)!}\cdot\zeta^{\prime}(1-4n),$
and $B_{2n}$ is the $2n$-th Bernoulli number. Moreover, as $s\to 0^{+}$, we have $g(s)=O(s^{2})$.
Proof.
The first part of the proposition follows from (10), (11), and the expansions, around $z=-4n$,
$s^{-z}=s^{4n}s^{-(z+4n)}=s^{4n}-(z+4n)\log s+O((z+4n)^{2}),$
and $\zeta(1+z)=\zeta(1-4n)+\zeta^{\prime}(1-4n)(z+4n)+O((z+4n)^{2})$ along with the fact that $\zeta(1-4n)=B_{2n}/4n$; see Titchmarch [13, p. 19].
For the second part, we start by using that
$\zeta^{\prime}(1-4n)=\frac{(-1)^{k+1}(4n)!}{2n(2\pi)^{2n}}\zeta^{\prime}(4n)+\frac{B_{4n}}{4n}\left(\sum_{k=1}^{4n-1}\frac{1}{k}-\gamma-\log(2\pi)\right).$
Using the fact that $|B_{4n}|\sim\tfrac{2(4n)!}{(2\pi)^{4n}}$ with Stirling’s approximation of the factorial, we have that
$\frac{b_{-4n}}{(4n)!}\cdot\zeta^{\prime}(1-4n)=O\left(\frac{(4n)!\log n}{(2\pi\sqrt{5})^{4n}n((2n)!)^{2}}\right)=O\left(\frac{\log n}{(2\pi\sqrt{5})^{4n}n^{3/2}}\right)=O(1).$
Now, we have at hand asymptotic information about all of the quantities in $\alpha_{n}$ except for $c_{-4n}$. It turns out that $c_{-4n}$ is uniformly bounded; more precisely $|c_{-4n}|<3$. To see this, note that for all $n\geqslant 1$, we have
$|c_{-4n}|=\left|\frac{1}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{\infty}{4n\choose k}\cdot\frac{(-1)^{k}}{\phi^{-4n+2k}+(-1)^{k+1}}\right|=\frac{1}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-2k}\left|\frac{1}{1-\phi^{4n-2k}}\right|$
since (${4n\choose k}=0$ for $k>4n$. Now the value of $\left|\tfrac{1}{1-\phi^{4n-2k}}\right|$ is maximal when $k=2n-1$ (recall, $k\neq 2n$), and there, it is approximately $2.6180$, which yields
$|c_{-4n}|<\frac{3}{5^{2n}}\sum_{\begin{subarray}{c}k=0\\ k\neq 2n\end{subarray}}^{4n}{4n\choose k}\cdot\varphi^{4n-2k}\leqslant\frac{3}{5^{2n}}\sum_{\begin{subarray}{c}k=0\end{subarray}}^{4n}{4n\choose k}\cdot\phi^{4n-k}\left(\frac{1}{\varphi}\right)^{k}=\frac{3}{5^{2n}}\left(\varphi+\frac{1}{\varphi}\right)^{4n}=3,$
which shows that $c_{-4n}$ is uniformly bounded. With this in hand, we will now use the fact that $|B_{2n}|=\tfrac{2(2n)!}{(2\pi)^{2n}}\zeta(2n)$ and $\zeta(2n)\sim 1$ as $n\to\infty$, to finish our proof. To this end, using the definition of $b_{-4n}$ and Stirling’s approximation, we have
$|\alpha_{n}|$ $<\frac{|B_{2n}|}{4n\cdot(4n)!}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$=\zeta(2n)\frac{2(2n)!}{4n\cdot(2\pi)^{2n}\cdot(4n)!}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$\sim\frac{2(2n)!}{4n\cdot(2\pi)^{2n}\cdot(4n)!}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{(4n)!}{((2n)!)^{2}}\log(4n)+3+\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+O(1)$
$=\frac{2}{4n\cdot(2\pi)^{2n}}\left(\frac{5^{-2n}}{\log\phi}\cdot\frac{\log(4n)}{(2n)!}+\frac{3\cdot(2n)!}{(4n)!}+\frac{5^{-2n}}{2(2n)!}\right)+O(1)$
$\sim\frac{2}{4n\cdot(2\pi)^{2n}}\cdot\frac{5^{-2n}}{\log\phi}\cdot\frac{\log(4n)}{(2n)!}+O(1)\sim\frac{1}{4\sqrt{\pi}\log{\varphi}}\cdot\frac{\log n}{\sqrt{n}}\cdot\left(\frac{e}{20\pi n}\right)^{2n}+O(1)=O(1).$
Similarly, we have that
$|\beta_{n}|\sim\frac{1}{4\sqrt{\pi}\log\varphi}\cdot\frac{1}{\sqrt{n}}\cdot\left(\frac{e}{20\pi n}\right)^{2n}=O(1).$
To see that $g(s)=O(s^{2})$ as $s\to 0^{+}$, we note that, using L’Hôpital’s rule, we have $\lim_{s\to 0^{+}}s^{2}\log s=0$. Using this, along with an application of the first part of the proposition,
$g(s)=\left(\alpha_{1}s^{4}-\beta_{1}s^{4}\log s\right)(1+O(s^{4}))=O(s^{2}).\qed$
Finally, at the simple poles of $\zeta_{F}(z)$ off the real line, which occur at $z=s(n,k)$ with $k\geqslant 0$ and $n\neq-k/2$, we note that $s^{-z}\Gamma(z)\zeta(1+z)$ is analytic at these points. Concerning the contributions from these singularities, we have the following result.
Proposition 5.
We have
$h(s):=$ $\,\sum_{k=0}^{\infty}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}\underset{{{z}={s(n,k)}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{F}(z)\right\}$
$=$ $\,\frac{1}{\log\varphi}\sum_{k=0}^{\infty}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}(-1)^{k}{-s(n,k)\choose k}\left(\frac{s}{\sqrt{5}}\right)^{-s(n,k)}\Gamma(s(n,k))\,\zeta(1+s(n,k)).$
Moreover, $h(s)=h(\varphi s)+O(s^{2})$ as $s\to 0^{+}$.
Proof.
The form $h(s)$ of these contributions is immediate using Proposition 1. Also, since the terms in each sum over $n$ is symmetric about the real axis, $h(s)$ is real. It remains to examine the analytic properties of $h(s)$ as a function of $s$.
We first consider the $k=0$ term of $h(s)$, which we denote by $[k=0]h(s)$. By a well-known properties of $\Gamma$ and $\zeta$, we have, for any real nonzero $t$, that, as $t\to\infty$, both $|\Gamma(it)|^{2}\sim\frac{\pi}{te^{\pi t}}$ and $\zeta(1+it)=O(\log|t|).$ Noting that for any complex number $z$, (${z\choose 0}=1$, we thus have that
$\big{|}[k=0]h(s)\big{|}\leqslant\sum_{n=1}^{\infty}\frac{O(\log n)}{\sqrt{n}\cdot e^{\frac{\pi^{2}n}{\log\varphi}}}=O(1).$
For $k\neq 0$, using the functional equations for $\zeta$ and $\Gamma$, since we are examining complex values, there is a positive constant $d_{n,k}$ that is uniformly bounded such that
$\big{|}\Gamma(s(n,k))\zeta(1+s(n,k))\big{|}$ $=\left|\frac{\zeta(-s(n,k))2^{s(n,k)}\pi^{s(n,k)+1}}{s(n,k)\sin\left(-\frac{s(n,k)\pi}{2}\right)}\right|$
$=\left|\frac{\pi\cdot\zeta\big{(}2k-i\tfrac{\pi(2n+k)}{\log\varphi}\big{)}}{\left(-2k+i\tfrac{\pi(2n+k)}{\log\varphi}\right)(2\pi)^{2k}\sin\left(i\tfrac{\pi^{2}(2n+k)}{2\log\varphi}\right)}\right|\sim d_{n,k}\cdot\frac{e^{-\frac{\pi^{2}k}{2\log\varphi}}}{(2\pi)^{2k}}\cdot e^{-\frac{\pi^{2}|n|}{\log\varphi}},$
where we have used that $|\sin(z)|$ is $\pi$-periodic in the $\Re(z)$, that $|e^{i\theta}|=1$ for all $\theta$, and that, as $k$ or $|n|$ (or both) grows, $\left|\sin\big{(}i\tfrac{\pi^{2}(2n+k)}{2\log\varphi}\big{)}\right|\sim{e^{\frac{\pi^{2}k}{2\log\varphi}}e^{\frac{\pi^{2}|n|}{\log\varphi}}}/{2}.$
It remains to deal with the factor (${-s(n,k)\choose k}$. To this end, note that
|($\left|{-s(n,k)\choose k}\right|=\left|\frac{\Gamma(1-s(n,k))}{k!\,\Gamma(1-s(n,k)-k)}\right|=\frac{1}{k!}\left(\prod_{j=k+1}^{2k}\left(j^{2}+\big{(}\tfrac{\pi(2n+k)}{\log\varphi}\big{)}^{2}\right)\right)^{1/2}\\ =\frac{(kn)^{k}}{k!}\left(\prod_{j=k+1}^{2k}\left(\left(\frac{j}{nk}\right)^{2}+\left(\tfrac{\pi\left(\frac{2}{k}+\frac{1}{n}\right)}{\log\varphi}\right)^{2}\right)\right)^{1/2}\leqslant\frac{(kn)^{k}}{k!}\left(\left(\frac{2}{n}\right)^{2}+\left(\tfrac{\pi\left(\frac{2}{k}+\frac{1}{n}\right)}{\log\varphi}\right)^{2}\right)^{k/2},$
and, independent of $k$,
$\left(\left(\frac{2}{n}\right)^{2}+\left(\tfrac{\pi\left(\frac{2}{k}+\frac{1}{n}\right)}{\log\varphi}\right)^{2}\right)^{k/2}\xrightarrow{n\to\infty}\left(\frac{2\pi}{k\log\varphi}\right)^{k}.$
Thus, there is a $d>0$, independent from $n$ and $k$, such that any term of $h(s)$ with $k\neq 0$, satisfies,
$\big{|}[k\neq 0]h(s)\big{|}$ $\leqslant d\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}\left(\frac{s}{\sqrt{5}}\right)^{2k}\frac{(kn)^{k}}{k!}\left(\frac{2\pi}{k\log\varphi}\right)^{k}\cdot d_{n,k}\cdot\frac{e^{-\frac{\pi^{2}k}{2\log\varphi}}}{(2\pi)^{2k}}\cdot e^{-\frac{\pi^{2}|n|}{\log\varphi}}$
$=d\left(\frac{s}{\sqrt{5}}\right)^{2k}\frac{1}{k!}\left(\frac{1}{\log\varphi}\right)^{k}\frac{e^{-\frac{\pi^{2}k}{2\log\varphi}}}{(2\pi)^{k}}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}d_{n,k}\cdot n^{k}\cdot e^{-\frac{\pi^{2}|n|}{\log\varphi}},$
where $d_{n,k}$ here is different from above, but still a uniformly bounded positive constant. So, there is a positive constant $d$, also different from above, but still independent of $n$ and $k$, such that
$\big{|}[k\neq 0]h(s)\big{|}$ $\leqslant d\left(\frac{s^{2}}{10\pi\cdot\log\varphi\cdot e^{\frac{\pi^{2}}{2\log\varphi}}}\right)^{k}\frac{1}{k!}\sum_{n=0}^{\infty}n^{k}e^{-\frac{\pi^{2}|n|}{\log\varphi}}$
$=d\left(\frac{s^{2}}{10\pi\cdot\log\varphi\cdot e^{\frac{\pi^{2}}{2\log\varphi}}}\right)^{k}\frac{1}{k!}\cdot\frac{e^{-\frac{\pi^{2}}{2\log\varphi}}A_{k}\big{(}e^{-\frac{\pi^{2}}{2\log\varphi}}\big{)}}{\big{(}1-e^{-\frac{\pi^{2}}{2\log\varphi}}\big{)}^{k+1}}$
$=d\left(\frac{s^{2}}{10\pi\cdot\log\varphi\cdot\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\right)^{k}\frac{1}{k!}\cdot\frac{A_{k}\big{(}e^{-\frac{\pi^{2}}{2\log\varphi}}\big{)}}{\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}$
(12) $<\frac{d}{\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\left(\frac{s^{2}}{10\pi\cdot\log\varphi\cdot\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\right)^{k},$
where, for $k\geqslant 1$, we have used that $\sum_{n\geqslant 0}n^{k}x^{n}={x\,A_{k}(x)}{(1-x)^{-(k+1)}},$ where $A_{k}(x)\in\mathbb{Z}_{>0}[x]$ is the $k$-th Eulerian Polynomial, which satisfies $\deg A_{k}(x)=k-1$ and $A_{k}(1)=k!$.
Finally, to see that $h(s)=h(\varphi s)+O(s^{2})$ as $s\to 0^{+}$, we note that by (12), the terms $[k\neq 0]h(s)$ contribute $O(s^{2})$ collectively, and for $[k=0]h(s)$, we note that
$\left(\frac{\varphi s}{\sqrt{5}}\right)^{\frac{i\pi(2n)}{\log\varphi}}$ $=\cos\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\varphi\alpha}{\sqrt{5}}\right)\right)+i\sin\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\varphi\alpha}{\sqrt{5}}\right)\right)$
$=\cos\left(2\pi n+\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)+i\sin\left(2\pi n+\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)$
$=\cos\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)+i\sin\left(\frac{2\pi n}{\log\varphi}\log\left(\frac{\alpha}{\sqrt{5}}\right)\right)=\left(\frac{s}{\sqrt{5}}\right)^{\frac{i\pi(2n)}{\log\varphi}}.\qed$
Lemma 2.
As $s\to 0^{+}$, we have
$\log(1-e^{-s})=\log s-2\gamma+O(s).$
Proof.
We follow the method above, writing
[…]
where, for now, $a>0$. We use the expansions of $\Gamma(z)$ and $\zeta(1+z)$ from above, to get that, as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma+O(s)$
which, when multiplied by $-1$, yields the desired result. ∎
Using the relationship in (2) and combining this lemma with the previous four propositions proves Theorem 3.
3. Fibonacci partitions $p_{F}(n)$ via the saddle point method
In this section, we prove our main result using a saddle point method. To achieve this, we must determine the behaviour as $s\to 0^{+}$ of each of the pieces in the expansion of $\log F(e^{-s})$.
We begin in the same way as Hardy and Ramanujan [4], using Cauchy’s integral formula, but diverge from their argument almost immediately. We have
[…]
where $C_{\varepsilon}$ indicates a positively oriented circle of radius $\varepsilon\in(0,1)$ and $s.p.$ indicates a path that goes through the saddle point $s=\alpha$ of the integrand, that is, the point $s=\alpha$ which is the solution of the equation
(13) $\frac{d}{ds}\left(s+\frac{1}{n}\log F_{2}(e^{-s})\right)=0.$
Our main result on Fibonacci partitions will follow from a result of Coons and Kirsten [3], which itself was inspired by the work of Nanda [7] and Richmond [10, 11].
Theorem 4 (Coons and Kirsten, 2009).
If $\Lambda(x)=\prod_{k\geqslant 1}\left(1\pm x^{\lambda_{k}}\right)^{\pm 1}$ generates a sequence $p_{\lambda}(n)$, and $s=\alpha$ is the solution of (13) with $F_{2}(e^{-s})$ replaced by $\Lambda(e^{-s})$, then, as $n$ tends to infinity,
$p_{\lambda}(n)=\frac{e^{n\alpha}\Lambda(e^{-\alpha})}{\sqrt{2\pi}}\left(\sqrt{\frac{1}{-\left.\frac{dn}{ds}\right|_{s=\alpha}}}+O\left(\frac{1}{n^{3/2}}\right)\right).$
Here, the $\pm$ indicates the same choice in both instances, and $\alpha$ must be thought of as being replaced by its large-$n$ asymptotic expansion so that the asymptotic of $p_{\lambda}(n)$ represents a large-$n$ asymptotic.
Remark 2.
The proof of Theorem 4 was accomplished by iteratively applying an asymptotic result on exponential integrals, which can be found in the book of Olver [9, Theorem 7.1]. Note that Coons and Kirsten [3] use the notation $t_{\Lambda}^{0}(n)$ for our definition of $p_{\lambda}(n)$ above. $\Diamond$
To apply this result, we must first determine the saddle point for large values of $n$. From (13), we have that
(14) $n=-\left.\frac{d}{ds}\log F_{2}(e^{-s})\right|_{s=\alpha}.$
We will use this combined with Theorem 3 to prove the following result.
Lemma 3.
For sufficiently large $n$, or, equivalently, for sufficiently small $\alpha>0$,
$n=\frac{-\log(\alpha\varphi e^{-c_{3}})}{\alpha\log\varphi}\left(1+O\left(\frac{\alpha}{\log\alpha}\right)\right).$
Proof.
We calculate
$n$ $=-\left.\frac{d}{ds}\log F_{2}(e^{-s})\right|_{s=\alpha}$
$=-\left.\frac{d}{ds}\left(\frac{(\log s)^{2}}{2\log\varphi}-(\log s)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(s)+O(s)\right)\right|_{s=\alpha}$
(15) $=-\frac{\log\alpha}{\alpha\log\varphi}+\frac{\log(e^{c_{3}}/\varphi)}{\alpha\log\varphi}+O(1),$
which, after factoring out the first two terms gives the result. ∎
Proposition 6.
For sufficiently small $\alpha>0$, or, equivalently, for sufficiently large $n$,
$\alpha=\frac{1}{n\log\varphi}\left(1+O\left(\frac{\log n}{n}\right)\right),$
and
$\left.\frac{dn}{ds}\right|_{s=\alpha}=-n^{2}(\log n)\log\varphi\left(1+O\left(\frac{1}{\log n}\right)\right).$
Proof.
It is quite clear that $O(\alpha/\log\alpha)=O(1/n).$ For the rest, we note that if $\log x=y\log y$, then for large $x$, $y=W(x)$, where $W$ is Lambert’s $W$-function, which satisfies
$W(x)=\log x(1+O(\log\log x/\log x))$
for all $x$ sufficiently large. Here, we have $x=\exp(n\log\varphi/\varphi e^{-c_{3}})$ and $y=1/\alpha\varphi e^{-c_{3}}$, so that
$\alpha\varphi e^{-c_{3}}=\frac{1}{W(\exp(n\log\varphi/\varphi e^{-c_{3}}))}\left(1+O\left(\frac{1}{n}\right)\right)=\frac{\varphi e^{-c_{3}}}{n\log\varphi}\left(1+O\left(\frac{\log n}{n}\right)\right),$
which proves the first assertion.
For the second assertion, we use (15) together with the first assertion to give
$\left.\frac{dn}{ds}\right|_{s=\alpha}=\frac{\log\alpha}{\alpha^{2}\log\varphi}\left(1+O\left(\frac{1}{\log\alpha}\right)\right)=-n^{2}(\log n)\log\varphi\left(1+O\left(\frac{1}{\log n}\right)\right).\qed$
We have the following immediate corollary of the first assertion of Proposition 6.
Corollary 1.
For sufficiently small $\alpha>0$, or, equivalently, for sufficiently large $n$,
$\log\alpha=-\log n-\log\log\varphi+O\left(\frac{\log n}{n}\right),$
and
$(\log\alpha)^{2}=(\log n)^{2}+2\log n\log\log\varphi+(\log\log\varphi)^{2}+O\left(\frac{(\log n)^{2}}{n}\right).$
Proposition 7.
Let $f(s):=g(s)+h(s)$, where $g(s)$ and $h(s)$ are as in Propositions 4 and 5, respectively. Then, as $s\to 0^{+}$, the function $f(s)$ satisfies $f(s)=f(\varphi s)+O(s^{2})$. That is, for sufficiently small $\alpha>0$, or equivalently, for sufficiently large $n$,
$f(\alpha)=\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n^{2}}\right),$
for some $1$-periodic function $\psi(x)$, and any $\varepsilon>0$.
Proof.
As $s\to 0^{+}$, using Proposition 4 and one application of L’Hôpital’s rule gives $g(s)=O(s^{2})$ (in fact, one gets that $g(s)=O(s^{4-\varepsilon})$ for any fixed small positive $\varepsilon$, but only $O(s^{2})$ is necessary after later comparison with the asymptotics for $h(s)$). For the result on $f(s)$, note that collectively, the sum of all of the terms with $k>0$ in the formula for $h(s)$ in Proposition 5 go to zero as $s\to 0^{+}$, since $\Re(-s(n,k))=2k>0$. Thus, as $s=\alpha\to 0^{+}$, separating out the $k=0$ term of the sum, and noting that $s(n,0)=\frac{2\pi in}{\log\varphi}$, we have,
(18) $f(\alpha)$ $=\frac{1}{\log\varphi}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq 0\end{subarray}}^{\infty}\left(\frac{\alpha}{\sqrt{5}}\right)^{-\frac{2\pi in}{\log\varphi}}\Gamma(\tfrac{2\pi in}{\log\varphi})\,\zeta(1+\tfrac{2\pi in}{\log\varphi})+O(\alpha^{2})$
$=\frac{2}{\log\varphi}\sum_{n=1}^{\infty}\Re\left(\left(\frac{\alpha}{\sqrt{5}}\right)^{-\frac{2\pi in}{\log\varphi}}\Gamma(\tfrac{2\pi in}{\log\varphi})\,\zeta(1+\tfrac{2\pi in}{\log\varphi})\right)+O(\alpha^{2})$
(19) $=\frac{2}{\log\varphi}\sum_{n=1}^{\infty}\Re\left(\left(\cos\left(\tfrac{2\pi n}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)\right)-i\sin\left(\tfrac{2\pi n}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)\right)\right)\right.$
×$\qquad\qquad\qquad\qquad\left.\times\Gamma(\tfrac{2\pi in}{\log\varphi})\,\zeta(1+\tfrac{2\pi in}{\log\varphi})\right)+O(\alpha^{2}).$
Since both the sine and cosine functions are $2\pi$-periodic, (19) gives that $f(s)=f(\varphi s)+O(s^{2})$ as $s\to 0^{+}$. Applying Proposition 6, we obtain
$\frac{2\pi j}{\log\varphi}\log\left(\tfrac{\alpha}{\sqrt{5}}\right)=-2\pi j\left(\frac{\log n}{{\log\varphi}}+\frac{\log\log\varphi+\log\sqrt{5}}{\log\varphi}\right)+O\left(\frac{\log n}{n}\right).$
Finally, we have $O(\alpha^{2})=O(n^{-2})$, which finishes the proof. ∎
We now have all of the elements to continue with our proof Theorem 1 in the case of non-distinct partitions of $n$.
Proof of Theorem 1 for $p_{F}(n)$.
We evaluate the pieces of the asymptotic in Theorem 4. Towards this end, the two parts of Proposition 6 give, respectively,
(20) $e^{n\alpha}=e^{n\cdot\frac{1}{n\log\varphi}\left(1+O\left(\frac{\log n}{n}\right)\right)}=e^{\frac{1}{\log\varphi}}\left(1+O\left(\frac{\log n}{n}\right)\right)$
and
(21) $\left(\sqrt{\frac{1}{-\left.\frac{dn}{ds}\right|_{s=\alpha}}}+O\left(\frac{1}{n^{3/2}}\right)\right)=\frac{1}{n}\left(\log n\log\varphi\right)^{-1/2}\left(1+O\left(\frac{1}{\log n}\right)\right).$
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $F_{2}(e^{-\alpha})$. To this end, we combine Theorem 3 with Proposition 7 to obtain, as $\alpha\to 0^{+}$,
$\log F_{2}(e^{-\alpha})$ $=\frac{(\log\alpha)^{2}}{2\log\varphi}-(\log\alpha)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+f(\alpha)+O(\alpha)$
$=\frac{1}{2\log\varphi}\left((\log n)^{2}+2\log n\log\log\varphi+(\log\log\varphi)^{2}+O\left(\frac{(\log n)^{2}}{n}\right)\right)$
$\quad+\left(\log n+\log\log\varphi+O\left(\frac{\log n}{n}\right)\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n}\right)$
$=\frac{(\log n)^{2}}{2\log\varphi}+(\log n)\left(\frac{\log\log\varphi}{\log\varphi}+\frac{c_{3}}{\log\varphi}-1\right)$
$\quad+\frac{(\log\log\varphi)^{2}}{2\log\varphi}+\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{(\log n)^{2}}{n}\right),$
where $\psi(x)$ is the $1$-periodic function described in Proposition 7. Thus, we have
$F_{2}(e^{-\alpha})=\psi_{1}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log n}{2\log\varphi}+\frac{\log\log\varphi}{\log\varphi}+\frac{c_{3}}{\log\varphi}-1\right)}\left(1+O\left(\frac{(\log n)^{2}}{n}\right)\right),$
where $\psi_{1}(x)$ is the $1$-periodic function satisfying
$\psi_{1}\left(\frac{\log n}{\log\varphi}\right):=\exp\left(\frac{(\log\log\varphi)^{2}}{2\log\varphi}+\left(\log\log\varphi\right)\left(\frac{c_{3}}{\log\varphi}-1\right)+c_{2}+2\gamma+\psi\left(\frac{\log n}{\log\varphi}\right)\right).$
Combing the asymptotic for $F_{2}(e^{-\alpha})$ with equations (20) and (21) gives the desired result, with
$\psi_{p_{F}}(x):=\frac{e^{\frac{1}{\log\varphi}}\psi_{1}(x)}{\sqrt{2\pi\log\varphi}}.\qed$
4. The normalised mean of distinct Fibonacci partitions: the function $q_{F}(n)$
In this section, we deal completely with the case of distinct partitions over Fibonacci numbers, that is, the case of $q_{F}(n)$ in Theorem 1. The same method of the previous two sections is employed, so for convenience and brevity, assuming the reader is acquainted with those sections, we move a bit quicker here. Towards this goal, we let $q_{F}(n)$ be the number of ways to write the integer $n$ as the sum of distinct Fibonacci numbers (with only one 1). We define
$G_{2}(x):=\sum_{n\geqslant 0}\left(\sum_{m\leqslant n}q_{F}(m)\right)x^{n}=(1-x^{2})^{-1}\prod_{k\geqslant 1}\left(1+x^{F_{k}}\right)$
and consider the function
$G(e^{-s}):=\prod_{k\geqslant 1}\left(1+x^{F_{k}}\right).$
Here, we have that
[…]
where $a>0$, $\eta(z)=\sum_{m\geqslant 1}\frac{(-1)^{m+1}}{m^{z}}=(1-2^{1-z})\zeta(z)$ denotes the Dirichlet $\eta$-function. Note that $\eta(z)$ is entire and shares the trivial zeros of $\zeta(z)$. Considering values, we have that $\eta(1)=\log(2)$ and $\eta(0)=1/2$.
Theorem 5.
The function $G(x)$ defined above satisfies, as $s\to 0^{+}$,
$\log G(e^{-s})=(-\log s)\frac{\log 2}{\log\varphi}+\frac{\log 2(4\gamma+2\log 5-\log\varphi-\log 2)}{2\log\varphi}+\frac{s}{2}+f_{1}(s),$
where $f_{1}(s)$ satisfies $f_{1}(s)=f_{1}(\varphi s)+O(s^{2})$.
Proof.
As above, combining our knowledge of the functions $s^{-z},$ $\Gamma(z)$ and $\eta(1+z)$ with Proposition 1, the integrand $s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)$ has
  • a double pole at $z=0$,
  • a simple pole coming from $\Gamma(z)$ at $z=-1$,
  • double poles at $z\in-4\mathbb{N}$, and
  • simple poles off the real line at $z=s(n,k)$, for $k\geqslant 0$ and $k\neq-2n$.
where, as before the poles of $\Gamma(z)$ at $z=-(4m+2)$ for $m\in\mathbb{N}_{0}$ are cancelled by the corresponding zeros of $\zeta_{F}(z)$ and the poles of $\Gamma(z)$ at odd negative integers $z=-(2m+1)$ for $m\in\mathbb{N}$ are cancelled by the corresponding zeros of $\eta(1+z)$.
Using (5) and (6), Proposition 1, Lemma 1, and
$\eta(1+z)=\log 2+\left(\gamma\log 2-\frac{(\log 2)^{2}}{2}\right)z+O(z^{2}),$
we see that the double pole at $z=0$ contributes
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}=(-\log s)\frac{\log 2}{\log\varphi}+(4\gamma+2\log 5-\log\varphi-\log 2)\frac{\log 2}{2\log\varphi}.$
The simple pole at $z=-1$ contributes
$\underset{{{z}={-1}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}=s\,\eta(0)\,\zeta_{F}(-1)\cdot\underset{{{z}={-1}}}{\textnormal{Res}}\left\{\Gamma(z)\right\}=s\left(\frac{1}{2}\right)(-1)(-1)=\frac{s}{2}.$
Using the proof of Proposition 4, replacing the contribution $\zeta(1-4n)=B_{2n}/4n$ with the contribution $\eta(1-4n)=(1-2^{4n})\zeta(1-4n)=(1-2^{4n})B_{2n}/4n$, we have
$g_{1}(s):=\sum_{n=1}^{\infty}\underset{{{z}={-4n}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}=\sum_{n=1}^{\infty}\alpha_{n}\,s^{4n}-\log s\sum_{n=1}^{\infty}\beta_{n}s^{4n},$
where $\beta_{n}:={b_{-4n}\eta(1-4n)}/{(4n)!}$,
$\alpha_{n}:=\frac{(1-2^{4n})B_{2n}}{4n\cdot(4n)!}\left(b_{-4n}\left(\sum_{k=1}^{4n}\frac{1}{k}-\gamma\right)+c_{-4n}-\frac{5^{-2n}}{2}\frac{(4n)!}{((2n)!)^{2}}\right)+\frac{b_{-4n}}{(4n)!}\cdot\eta^{\prime}(1-4n),$
and $B_{2n}$ is the $2n$-th Bernoulli number. Moreover, as $s\to 0^{+}$, we have $g_{1}(s)=O(s^{2})$.
Finally, we have
$h_{1}(s):=$ $\,\sum_{k=0}^{\infty}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}\underset{{{z}={s(n,k)}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{F}(z)\right\}$
$=$ $\,\frac{1}{\log\varphi}\sum_{k=0}^{\infty}\sum_{\begin{subarray}{c}n=-\infty\\ n\neq-k/2\end{subarray}}^{\infty}(-1)^{k}{-s(n,k)\choose k}\left(\frac{s}{\sqrt{5}}\right)^{-s(n,k)}\Gamma(s(n,k))\,\eta(1+s(n,k)).$
Now, since $|(1-2^{it})|<2,$ for all real $t$, we have, from the result for $\zeta(1+it)$ that $\eta(1+it)=O(\log|t|)$, so that $\big{|}[k=0]h_{1}(s)\big{|}=O(1)$. For $k\neq 0$, we note that
$\big{|}\Gamma(s(n,k))\eta(1+s(n,k))\big{|}=\big{|}1-2^{-s(n,k)}\big{|}\cdot\big{|}\Gamma(s(n,k))\zeta(1+s(n,k))\big{|}<16^{k}\big{|}\Gamma(s(n,k))\zeta(1+s(n,k))\big{|}.$
Thus, in the proof of Proposition 5, we have the analogue of (12), with the added $16^{k}$ factor,
$\big{|}[k\neq 0]h_{1}(s)\big{|}<\frac{d}{\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\left(\frac{16s^{2}}{10\pi\cdot\log\varphi\cdot\big{(}e^{\frac{\pi^{2}}{2\log\varphi}}-1\big{)}}\right)^{k},$
which completes the proof of the theorem. The proof that $f_{1}(s):=g_{1}(s)+h_{1}(s)=f_{1}(\varphi s)+O(s^{2})$ follows mutatis mutandis the proof of Proposition 7. ∎
Corollary 2.
The function $G_{2}(x)$ defined above satisfies, as $s\to 0^{+}$,
$\log G_{2}(e^{-s})=(-\log s)\left(\frac{\log 2}{\log\varphi}+1\right)+\frac{\log 2(4\gamma+2\log 5-\log\varphi-\log 2)}{2\log\varphi}+2\gamma-\log 2+f_{1}(s)+O(s),$
where $f_{1}(s)$ satisfies $f_{1}(s)=f_{1}(\varphi s)+O(s^{2})$.
Proof.
We need only determine the asymptotics of $-\log(1-e^{-2s})$ and add those to the result in Theorem 5. To this end, note that $1-e^{-2s}=(1+e^{-s})(1-e^{-s})$. Now, as $s\to 0^{+}$, we have that $-\log(1+e^{-s})=-\log 2+O(s).$ To find the asymptotics of $-\log(1-e^{-s})$, we appeal to Lemma 2, which gives as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma+O(s),$
so that $-\log(1-e^{-2s})=-\log s+2\gamma-\log 2+O(s),$ which when added to the statement in Theorem 5 yields the desired result. ∎
We find the saddle point as in the previous section. Here, we note that the saddle point $s=\alpha$ is the solution to
$\frac{d}{ds}\left(s+\frac{1}{n}\log G_{2}(e^{-s})\right)=0.$
Here, we use the asymptotics in Corollary 2 to give that
$n=-\left.\frac{d}{ds}\log G_{2}(e^{-s})\right|_{s=\alpha}=\frac{1}{\alpha}\left(\frac{\log 2}{\log\varphi}+1\right)+O(1)=\frac{1}{\alpha}\left(\frac{\log 2}{\log\varphi}+1\right)\left(1+O(\alpha)\right).$
Thus,
$\left.\frac{dn}{ds}\right|_{s=\alpha}=-\frac{1}{\alpha^{2}}\left(\frac{\log 2}{\log\varphi}+1\right)\left(1+O(\alpha^{2})\right).$
This gives the following result.
Proposition 8.
For sufficiently small $\alpha>0$, or, equivalently, for sufficiently large $n$,
$\alpha=\frac{1}{n}\left(\frac{\log 2}{\log\varphi}+1\right)\left(1+O\left(\frac{1}{n}\right)\right),$
and
$\left.\frac{dn}{ds}\right|_{s=\alpha}=-\frac{n^{2}}{\left(\frac{\log 2}{\log\varphi}+1\right)}\left(1+O\left(\frac{1}{n^{2}}\right)\right).$
Corollary 3.
For sufficiently small $\alpha>0$, or equivalently, for sufficiently large $n$,
$-\log\alpha=\log n+\log\left(\frac{\log 2}{\log\varphi}+1\right)+O\left(\frac{1}{n}\right).$
We are now prepared to finish the proof of Theorem 1. Again, we apply Theorem 4.
Proof of Theorem 1 for $q_{F}(n)$.
As before, we evaluate the necessary pieces Theorem 4, with the appropriate generating function and saddle points. Proposition 8 gives, for any fixed $\varepsilon>0$,
(22) $e^{n\alpha}=e^{\left(\frac{\log 2}{\log\varphi}+1\right)}\left(1+O\left(\frac{1}{n}\right)\right)$
and
(23) $\left(\sqrt{\frac{1}{-\left.\frac{dn}{da}\right|_{a=\alpha}}}+O\left(\frac{1}{n^{3/2}}\right)\right)=\frac{1}{n}\sqrt{\frac{\log 2}{\log\varphi}+1}\left(1+O\left(\frac{1}{n^{2}}\right)\right).$
It remains to determine the small $\alpha$, and hence large $n$, asymptotics of $G_{2}(e^{-\alpha})$. To this end, we combine Corollary 2 with Corollary 3 to obtain, as $\alpha\to 0^{+}$, or, equivalently, as $n\to\infty$,
$\log G_{2}(e^{-\alpha})$ $=(\log n)\left(\frac{\log 2}{\log\varphi}+1\right)+\left(\frac{\log 2}{\log\varphi}+1\right)\log\left(\frac{\log 2}{\log\varphi}+1\right)$
$\qquad\qquad+\frac{\log 2(4\gamma+2\log 5-\log\varphi-\log 2)}{2\log\varphi}+2\gamma-\log 2+\psi_{2}\left(\frac{\log n}{\log\varphi}\right)+O\left(\frac{1}{n}\right),$
where $\psi_{2}(x)$ is an explicit $1$-periodic function. Thus, we have
$G_{2}(e^{-\alpha})=\psi_{3}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log 2}{\log\varphi}+1\right)}\left(1+O\left(\frac{1}{n}\right)\right),$
where $\psi_{3}(x)$ is an explicit $1$-periodic function. Combining this with equations (22) and (23) finishes the proof of the theorem. ∎
5. Partitions over general linear recurrences with a dominant root
The results of the previous sections on Fibonacci partitions can be generalised to positive recurrence sequences $P_{n}$ with an irreducible characteristic polynomial having a positive dominant real root and such that $P_{1}=1$. We will also assume that the values $P_{n}$ are distinct. Here, we require $P_{1}=1$ so that there always exists a non-distinct partition of $n$ over $P_{n}$. The restriction on the values of $P_{n}$ being distinct is not a very strict one—since the recursion has a dominant real root there are at most finitely many repeated values, so an analysis analogous to moving between $F(x)$ and $F_{2}(x)$ (or, $G(x)$ and $G_{2}(x)$) is possible, and uncomplicated. For such $P_{n}$, we wish to asymptotically understand the number of solutions to
(24) $n=a_{1}P_{1}+a_{2}P_{2}+\cdots+a_{k}P_{K}+\cdots.$
As before, we let $p_{P}(n)$ denote the number of non-distinct partitions of $n$ over the sequence $P_{n}$, that is, solutions to (1) in nonnegative integers $a_{k}$, and let $q_{P}(n)$ denote the number of distinct partitions of $n$ over the sequence $P_{n}$, that is, solutions to (1) with $a_{k}\in\{0,1\}$.
The case of partitions over $P_{n}$ is carried out exactly as in the case with the Fibonacci numbers $F_{n}$, but with the Fibonacci zeta function replaced by the zeta function $\zeta_{P}(z)$ which is the meromorphic continuation of the Dirichlet series $\sum_{k\geqslant 1}P_{k}^{-z}$ ($\Re(z)>0$). Here, we consider the generating functions,
$F_{P}(x):=\sum_{n\geqslant 0}p_{P}(n)x^{n}=\prod_{k\geqslant 1}\left(1-x^{P_{k}}\right)^{-1}$
and
$G_{P}(x):=\sum_{n\geqslant 0}\left(\sum_{m\leqslant n}q_{P}(m)\right)x^{n}=(1-x)^{-1}\prod_{k\geqslant 1}\left(1+x^{P_{k}}\right).$
To complete our analysis, we use the following result of Serrano Holdago and Navas Vicente [12], which is a generalisation of Navas [8].
Proposition 9 (Serrano Holdago and Navas Vicente, 2023).
Let $P(x)$ be the minimal polynomial with $\deg P(x)=r$ of the linear recurrence $P_{n}$ (as described above) and let $\beta>1$ be the dominant root of $P(x)$. Then the Dirichlet series $\zeta_{P}(z):=\sum_{k\geqslant 1}P_{k}^{-z}$ ($\Re(z)>0$) can be analytically continued to a meromorphic function, also denoted $\zeta_{P}(z)$, all of whose singularities are simple poles at the points
$s{(n,\boldsymbol{k})}:=\frac{\log|\beta^{-k_{1}}\beta_{2}^{k_{1}-k_{2}}\cdots\beta_{r}^{k_{r-1}}|}{\log\beta}+i\cdot\frac{\arg(\beta_{1}^{-k_{1}}\beta_{2}^{k_{1}-k_{2}}\cdots\beta_{r}^{k_{r-1}})+2\pi n}{\log\beta},$
where $\boldsymbol{k}=(k_{1},\ldots,k_{r-1})$, $\beta_{2},\ldots,\beta_{r}$ are the algebraic conjugates of $\beta$, $n\in\mathbb{Z}$ and the parameters $k_{1},\ldots,k_{r-1}$ are integers satisfying $0\leqslant k_{r-1}\leqslant k_{r-2}\leqslant\cdots\leqslant k_{1}$.
We adopt the terminology of Proposition 9 for the rest of this section along with the definitions of the real numbers $\lambda=\lambda_{1},\lambda_{2},\ldots,\lambda_{r}$ satisfying
$P_{n}=\lambda\beta^{n}+\lambda_{2}\beta_{2}^{n}+\cdots+\lambda_{r}\beta_{r}^{n}.$
Note that since $P_{n}$ is strictly increasing, we necessarily have that $\lambda>0$.
Continuing the analogy with the Fibonacci partitions, we need asymptotic results, as $s\to 0^{+}$, of the functions
[…]
As before, we have a few different types of poles to consider:
  • a triple pole (resp., double pole) at $z=0$ in the case of $F_{P}(e^{-s})$ (resp., $G_{P}(e^{-s})$),
  • simple poles at countable (and separated) non-integer real values $z\leqslant-\frac{\log(\beta/|\beta_{2}|)}{\log\beta}.$
  • double poles at $z\in-\mathbb{N}$, and
  • simple poles off the real line at $z=s(n,\boldsymbol{k})$.
Now, it is clear from the previous analysis on Fibonacci partitions that the main contribution will come from the pole at $z=0$, which comes only from $\boldsymbol{k}=\boldsymbol{0}=(0,\ldots,0).$ The negative real poles will give a cumulative contribution of $O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for any fixed $\varepsilon>0$, since $s\log s=O(s^{1-\varepsilon})$ for any fixed $\varepsilon>0$ as $s\to 0^{+}$. As well, the simple poles off the real line contribute a function $f_{2}(s)$ (resp., $f_{3}(s)$) toward $\log F_{P}(e^{-s})$ (resp., $\log G_{P}(e^{-s})$), that satisfy $f_{i}(s)=f_{i}(\beta s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for $i\in\{2,3\}$. It remains to obtain the contributions from the pole at $z=0$. For these, we require the following result.
Lemma 4.
Near $z=0$, we have
$\zeta_{P}(z)=\frac{1}{\log\beta}\cdot\frac{1}{z}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right)z+O(z^{2}),$
where $C_{1}:=\sum_{k\geqslant 1}\sum_{\boldsymbol{k}\neq\boldsymbol{0}}\frac{1}{k_{1}}{k_{1}\choose k_{2}}\cdots{k_{r-2}\choose k_{r-1}}\frac{(\lambda_{r}\beta_{r}^{k})^{k_{r-1}}}{(\lambda\beta^{k})^{z+k_{1}}}\left(\prod_{j=2}^{r-1}(\lambda_{j}\beta_{j}^{k})^{k_{j+1}-k_{j}}\right)$.
Proof.
Following Serrano Holdago and Navas Vicente [12], we write
$\zeta_{P}(z)$ $=\sum_{k\geqslant 1}\sum_{\boldsymbol{k}}{-z\choose k_{1}}{k_{1}\choose k_{2}}\cdots{k_{r-2}\choose k_{r-1}}\frac{(\lambda_{r}\beta_{r}^{k})^{k_{r-1}}}{(\lambda\beta^{k})^{z+k_{1}}}\left(\prod_{j=2}^{r-1}(\lambda_{j}\beta_{j}^{k})^{k_{j+1}-k_{j}}\right)$
$=\sum_{k\geqslant 1}\frac{1}{(\lambda\beta^{k})^{z}}+\sum_{k\geqslant 1}\sum_{\boldsymbol{k}\neq\boldsymbol{0}}{-z\choose k_{1}}{k_{1}\choose k_{2}}\cdots{k_{r-2}\choose k_{r-1}}\frac{(\lambda_{r}\beta_{r}^{k})^{k_{r-1}}}{(\lambda\beta^{k})^{z+k_{1}}}\left(\prod_{j=2}^{r-1}(\lambda_{j}\beta_{j}^{k})^{k_{j+1}-k_{j}}\right)$
$=\frac{\lambda^{-z}}{\beta^{z}-1}+\sum_{k\geqslant 1}\sum_{\boldsymbol{k}\neq\boldsymbol{0}}\frac{\Gamma(z+k_{1})}{\Gamma(z)\,k_{1}!}{k_{1}\choose k_{2}}\cdots{k_{r-2}\choose k_{r-1}}\frac{(\lambda_{r}\beta_{r}^{k})^{k_{r-1}}}{(\lambda\beta^{k})^{z+k_{1}}}\left(\prod_{j=2}^{r-1}(\lambda_{j}\beta_{j}^{k})^{k_{j+1}-k_{j}}\right).$
where $\sum_{\boldsymbol{k}}:=\sum_{k_{1}\geqslant 0}\sum_{k_{2}=0}^{k_{1}}\cdots\sum_{k_{r-1}=0}^{k_{r-2}}.$ Again, using that $\tfrac{1}{\Gamma(z)}=z+O(z^{2})$ and $\Gamma(z+k_{1})=(k_{1}-1)!+O(z)$ near $z=0$, along with the asymptotic expansions
$\lambda^{-z}=1-z\log\lambda+\frac{1}{2}(\log\lambda)^{2}z^{2}+O(z^{3}),$
and
$\frac{1}{\beta^{z}-1}=\frac{1}{\log\beta}\cdot\frac{1}{z}-\frac{1}{2}+\frac{\log\beta}{12}z+O(z^{2}),$
we have that
$\zeta_{P}(z)=\frac{1}{\log\beta}\cdot\frac{1}{z}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right)z+O(z^{2}),$
where $C_{1}:=\sum_{k\geqslant 1}\sum_{\boldsymbol{k}\neq\boldsymbol{0}}\frac{1}{k_{1}}{k_{1}\choose k_{2}}\cdots{k_{r-2}\choose k_{r-1}}\frac{(\lambda_{r}\beta_{r}^{k})^{k_{r-1}}}{(\lambda\beta^{k})^{z+k_{1}}}\left(\prod_{j=2}^{r-1}(\lambda_{j}\beta_{j}^{k})^{k_{j+1}-k_{j}}\right)$.∎
In the following result, we use Lemma 4 to determine the asymptotic behaviour of the functions $\log F_{P}(e^{-s})$ and $\log G_{P}(e^{-s})$ as $s\to 0^{+}$. All of the expansions of the functions involved, $s^{-z}$, $\Gamma(z)$, $\zeta(1+z)$ and $\eta(1+z)$, have been noted somewhere in the previous sections of this work—we use them below without further reference.
Proposition 10.
For any fixed $\varepsilon\in(0,1)$, as $s\to 0^{+}$, the functions $F_{P}(z)$ and $G_{P}(z)$ defined above satisfy,
$\log F_{P}(e^{-s})=\frac{(\log s)^{2}}{2\log\beta}+\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}-\frac{2\gamma}{\log\beta}\right)\log s+C_{2}+f_{2}(s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}),$
and
$\log G_{P}(e^{-s})=-\left(\frac{\log 2}{\log\beta}+1\right)\log s+C_{4}+f_{3}(s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}),$
where
$C_{2}:=\frac{1}{\log\beta}\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}\right)-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)2\gamma+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right),$
$C_{4}:=2\gamma+C_{3}=2\gamma-\log 2\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)+\frac{1}{\log\beta}\left(2\gamma-\frac{(\log 2)^{2}}{2}\right),$
and $f_{i}(s)=f_{i}(\beta s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for $i\in\{2,3\}$.
Proof.
For $F_{P}(e^{-s})$, note that near $z=0$, we have
$s^{-z}\Gamma(z)\zeta(1+z)=\frac{1}{z^{2}}+\left(2\gamma-\log s\right)\frac{1}{z}+\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}-2\gamma\log s+\frac{(\log s)^{2}}{2}\right)+O(z),$
so that, using Lemma 4, we have that
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\zeta(1+z)\zeta_{P}(z)\right\}$ $=\frac{1}{\log\beta}\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}-2\gamma\log s+\frac{(\log s)^{2}}{2}\right)$
$\qquad-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)(2\gamma-\log s)+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right)$
$=\frac{(\log s)^{2}}{2\log\beta}+\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}-\frac{2\gamma}{\log\beta}\right)\log s+C_{2},$
where $C_{2}:=\frac{1}{\log\beta}\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}\right)-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)2\gamma+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right).$ The contributions coming from the rest of the poles are as discussed before the statement of Lemma 4.
Again, near $z=0$, considering now the case $G_{P}(e^{-s})$, we have
$s^{-z}\Gamma(z)\eta(1+z)=\frac{\log 2}{z}+\left(2\gamma-\frac{(\log 2)^{2}}{2}\right)-\log 2\log s+O(z),$
so that, using Lemma 4, we have that
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{P}(z)\right\}$ $=-\frac{\log 2}{\log\beta}\log s+C_{3},$
where
$C_{3}:=-\log 2\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)+\frac{1}{\log\beta}\left(2\gamma-\frac{(\log 2)^{2}}{2}\right).$
We also need to add in the contribution from $-\log(1-e^{-s})$, which using the expansions from Section 2, is, as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma+O(s).$
As in the above case of $F_{P}(e^{s})$, the contributions coming from the rest of the poles are as discussed before the statement of Lemma 4. ∎
On inspection, one notices that the dominant asymptotic terms of both $\log F_{P}(e^{-s})$ and $\log G_{P}(e^{-s})$ are precisely the dominant asymptotic terms of $\log F_{2}(e^{-s})$ and $\log G_{2}(e^{-s})$ of the functions related to the Fibonacci partitions, after substituting $\beta$ for $\varphi$. Of course, this is not so surprising, as the Fibonacci numbers $F_{k}$ are just a special case of the more general sequence $P_{k}$. The property of note, here, is regarding the associated saddle points. Since the leading order behaviour is the same, the saddle points satisfy the same leading order asymptotics. In particular, Proposition 6 and Corollary 1 hold for the saddle point $\alpha_{F_{P}}$ related to non-distinct partitions of $n$ over $P_{k}$, and so also then, do (20) and (21). Similarly, Proposition 8, Corollary 3, (22) and (23), hold for the saddle point $\alpha_{G_{P}}$ related to the partial sums of distinct partitions of $n$ over $P_{k}$. We use these results, as well as this notation, below.
Proof of Theorem 2.
We first consider the case of $p_{P}(n)$, so, of $F_{P}(e^{-\alpha_{F_{p}}})$. Then, by Proposition 10, Proposition 6 and Corollary 1, the small-$\alpha_{F_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
$\log F_{P}(e^{-\alpha_{F_{p}}})$ $=\frac{(\log\alpha_{F_{p}})^{2}}{2\log\beta}-(\log\alpha_{F_{p}})\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad+C_{2}+f_{2}(\alpha_{F_{p}})+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$
$=\frac{1}{2\log\beta}\left((\log n)^{2}+2\log n\log\log\beta+(\log\log\beta)^{2}+O\left(\frac{(\log n)^{2}}{n}\right)\right)$
$\qquad\qquad+\left(\log n+\log\log\beta+O\left(\frac{\log n}{n}\right)\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)$
$=\frac{(\log n)^{2}}{2\log\beta}+(\log n)\left(\frac{\log\log\beta}{\log\beta}+\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)\right)$
$\qquad\qquad+\frac{(\log\log\beta)^{2}}{2\log\beta}+\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right),$
where $\psi_{4}(x)$ is an explicitly determinable $1$-periodic function. Thus, we have
$F_{P}(e^{-\alpha_{F_{P}}})=\psi_{5}\left(\frac{\log n}{\log\beta}\right)n^{\left(\frac{\log n}{2\log\beta}+\frac{\log\log\beta}{\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)}\left(1+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)\right),$
where $\psi_{5}(x)$ is the $1$-periodic function satisfying
$\psi_{5}\left(\frac{\log n}{\log\beta}\right):=\exp\left(\frac{(\log\log\beta)^{2}}{2\log\beta}+\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)\right).$
Combing the asymptotic for $F_{P}(e^{-\alpha_{F_{P}}})$ with equations (20) and (21) gives the desired result for $p_{P}(n)$, with $\psi_{p_{P}}(x):={e^{\frac{1}{\log\beta}}\psi_{1}(x)}/{\sqrt{2\pi\log\varphi}}.$
We now consider the number $q_{P}(n)$ of distinct partitions over $P_{k}$, and so the asymptotic behaviour of $G_{P}(e^{-\alpha_{G_{p}}})$. Then, by Proposition 10, Proposition 8 and Corollary 3, the small-$\alpha_{G_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
$\log G_{P}(e^{-\alpha_{G_{p}}})$ $=(\log n)\left(\frac{\log 2}{\log\beta}+1\right)+\left(\frac{\log 2}{\log\beta}+1\right)\log\left(\frac{\log 2}{\log\beta}+1\right)$
$\qquad\qquad+C_{4}+\psi_{6}\left(\frac{\log n}{\log\beta}\right)+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right),$
where $\psi_{6}(x)$ is an explicit $1$-periodic function. Thus, we have
$G_{P}(e^{-\alpha_{G_{p}}})=\psi_{7}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log 2}{\log\beta}+1\right)}\left(1+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)\right),$
where $\psi_{7}(x)$ is an explicit $1$-periodic function. The asymptotic for $G_{P}(e^{-\alpha})$ combined with the contributions from equations (22) and (23) finishes the proof of the theorem. ∎
Acknowledgements
This work was started during a short visit of M. Coons to Aarhus University; he thanks their faculty and staff for their hospitality and support. S. Kristensen and M. L. Laursen are supported by the Independent Research Fund Denmark (Grant ref. 1026-00081B) and the Aarhus University Research Research Foundation (Grant ref. AUFF-E-2021-9-20). The idea for this work came about when M. Coons was a master’s student working under Klaus Kirsten, and certainly this work benefitted from several conversations with Klaus and also Markus Hunziker, which occurred nearly twenty years ago! M. Coons gives a very belated thanks to both of them.